# Statistical Inference for Policy Evaluation with Temporal Difference Learning

### v3

@@ B.1.3 The necessity of the condition (71)
An important fact to note about Theorem 3.1 is that it relies on choosing an appropriate initial stepsize $\eta_{0}$ dependent of $\delta$. The relationship is specified in Equation (71). This constraint, albeit only related to $\delta$ in logarithm terms, prevents $\delta$ to converge to 00 for any fixed $\eta_{0}$. One might naturally wonder if such a constraint could be relaxed and the convergence of $\overline{\bm{\Delta}}_{T}$ can be guaranteed with high probability for any fixed $\eta_{0}$ and arbitrarily small $\delta$. The following proposition shows that the answer is negative, so some form of constraint on $\delta$ is necessary for a Chernoff-style convergence guarantee.
@@
[…]
@@
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\tilde{C}\;T^{-\frac{1}{2}},$
where $\tilde{C}$ is a problem related quantity independent of $T$.
@@
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\tilde{C}^{\prime}\;T^{2\alpha-2},$
where $\tilde{C}^{\prime}$ is a problem related quantity independent of $T$.
@@
Defining $\bm{W}$,$\bm{D}$ and other statistics.
@@
since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$. This effectively means that the last term on the right-hand-side of (B.2.1) is 00. Therefore, the difference bewteen $\Delta$ and $\Delta^{(i)}$ is bounded by
@@
$\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ =$=\begin{cases}(\bm{A}_{i}-\check{\bm{A}}_{i})(\bm{\theta}_{i-1}-\bm{\theta}^{\star}),\quad\text{for}\quad j=i;\\ (\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right),\quad\text{for}\quad j>i.\end{cases}$
@@
11 11 $=\eta_{0}\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}\bm{Y}+\bm{Y}\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
@@
$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|$ |($\leq\underset{I_{1}}{\underbrace{\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}$
+|($\quad+\underset{I_{2}}{\underbrace{\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}+\underset{I_{3}}{\underbrace{\left|\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}.$
@@ Appendix C Proof of supportive lemmas and propositions
$\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ $=\underset{I_{1}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}}$
$+\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}]-\mathbb{E}[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}},$
@@
$I_{1}$ $=\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]$
$=\underset{I_{11}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]}}+\underset{I_{12}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]}}.$
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
$\leq\frac{1}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right)}$
$\leq\frac{1}{T}\sum_{t=1}^{T}4\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}=4\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}.$ (107)
@@
$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\tilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
the Markov’s inequality indicates
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}\leq\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}+o(T^{-\frac{1}{2}})$ (108)
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ (109)
$\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
$\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
$\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
$\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
$=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (110)
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$
$\leq 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (111)
@@
Hence by combining (108) and (C.4), we obtain
@@
$\left\|I_{2}\right\|_{\mathsf{F}}\leq\frac{2\sqrt{2}}{\sqrt{T}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\sqrt{\log\frac{3}{\delta}}$ (113)
with probability at least $1-\delta$. Lemma B.8 follows by combining (112) and (113) with $\delta=T^{-\frac{1}{3}}$.∎

### v4

@@ B.1.3 The necessity of the condition (74)
An important fact to note about Theorem 3.1 is that it relies on choosing an appropriate initial stepsize $\eta_{0}$ dependent of $\delta$. The relationship is specified in Equation (74). This constraint, albeit only related to $\delta$ in logarithm terms, prevents $\delta$ to converge to 0 for any fixed $\eta_{0}$. One might naturally wonder if such a constraint could be relaxed and the convergence of $\overline{\bm{\Delta}}_{T}$ can be guaranteed with high probability for any fixed $\eta_{0}$ and arbitrarily small $\delta$. The following proposition shows that the answer is negative, so some form of constraint on $\delta$ is necessary for a Chernoff-style convergence guarantee.
@@
It is easy to verify that $\bm{\mu}=(0.1,0.1,0.8)^{\top}$. Furthermore, let d=1d=1333In this case, many vectors and matrices reduce to scalars; however, we keep the boldface notation for simplicity., $\bm{\phi}(1)=-0.5$$\bm{\phi}(2)=-1$ and $\bm{\phi}(3)=1$. Direct calculation then yields
@@
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\widetilde{C}\;T^{-\frac{1}{2}},$
where $\widetilde{C}$ is a problem related quantity independent of $T$.
@@
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}^{\prime}\;T^{2\alpha-2},$
where $\widetilde{C}^{\prime}$ is a problem related quantity independent of $T$.
@@
Defining $\bm{W}$$\bm{D}$ and other statistics.
@@
since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$. This effectively means that the last term on the right-hand-side of (B.2.1) is 0. Therefore, the difference between $\Delta$ and $\Delta^{(i)}$ is bounded by
@@
$\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ $=\begin{cases}(\bm{A}_{i}-\check{\bm{A}}_{i})(\bm{\theta}_{i-1}-\bm{\theta}^{\star}),\quad\text{for}\quad j=i;\\ (\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right),\quad\text{for}\quad j>i.\end{cases}$
@@
$1$ $=\eta_{0}\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}\bm{Y}+\bm{Y}\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
@@
$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|$ $\leq\underset{I_{1}}{\underbrace{\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}$
$\quad+\underset{I_{2}}{\underbrace{\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}+\underset{I_{3}}{\underbrace{\left|\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}.$
@@ Appendix C PROOF OF SUPPORTIVE LEMMAS AND PROPOSITIONS
$\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ $=\underset{I_{1}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}}$
$+\underset{I_{2}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}]-\mathbb{E}[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}},$
@@
$I_{1}$ $=\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]$
$=\underset{I_{11}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]}}+\underset{I_{12}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]}}.$
@@
$\left\|\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq\sum_{t=1}^{T}w_{t,T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right\|_{\mathsf{F}}$
$\leq\sum_{t=1}^{T}w_{t,T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right)}$
$\leq\sum_{t=1}^{T}w_{t,T}\cdot 4\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}.$ (110)
@@
$\mathbb{E}\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}$ $=\frac{1}{t}\mathsf{Tr}(\overline{\bm{\Lambda}}_{t})+\widetilde{C}t^{-\alpha-1}=\frac{1}{t}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(t^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}t}+o(t^{-1}),$
the expectation of this bound can be bounded by
$\mathbb{E}\left\|\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq 4\sum_{t=1}^{T}w_{t,T}\cdot\mathbb{E}\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}$
$\leq 4\sum_{t=1}^{T}\frac{2t}{T(T+1)}\cdot\left(\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}t}+o(t^{-1})\right)$
$\asymp\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1});$
Consequently, the Markov’s inequality indicates
$\left\|I_{11}\right\|_{\mathsf{F}}\leq\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}+o(T^{-\frac{1}{2}})$ (111)
@@
$\left\|\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]\right\|_{\mathsf{F}}$
$\leq\sum_{t=1}^{T}w_{t,T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right\|_{\mathsf{F}}$
$\leq 2\sum_{t=1}^{T}w_{t,T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
$\leq 2\sum_{t=1}^{T}w_{t,T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
$\leq 4\sum_{t=1}^{T}w_{t,T}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{t}\|_{2}$
$=4(2\|\bm{\theta}^{\star}\|_{2}+1)\sum_{t=1}^{T}w_{t,T}\|\overline{\bm{\Delta}}_{t}\|_{2}.$ (112)
Furthermore, since $\sum_{t=1}^{T}w_{t,T}=1$, the Jensen’s inequality guarantees
$\mathbb{E}\left(\sum_{t=1}^{T}w_{t,T}\|\overline{\bm{\Delta}}_{t}\|_{2}\right)^{2}$ $\leq\sum_{t=1}^{T}w_{t,T}\mathbb{E}\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}$
$\leq\sum_{t=1}^{T}\frac{2t}{T(T+1)}\left(\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}t}+o(t^{-1})\right)$
$\asymp\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1})$
@@
$\left\|I_{22}\right\|_{\mathsf{F}}\lesssim 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (113)
@@
Hence by combining (111) and (113), we obtain
@@
$\left\|I_{2}\right\|_{\mathsf{F}}\leq 2\sqrt{2}\sqrt{\sum_{t=1}^{T}w_{t,T}^{2}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\sqrt{\log\frac{3}{\delta}};$ (115)
with probability at least $1-\delta$. Since
$\sum_{t=1}^{T}w_{t,T}^{2}$ $=\frac{4}{T^{2}(T+1)^{2}}\sum_{t=1}^{T}t^{2}$
$=\frac{4}{T^{2}(T+1)^{2}}\cdot\frac{T(T+1)(2T+1)}{6}\asymp\frac{1}{T},$
Lemma B.8 follows by combining (114) and (115) with $\delta=T^{-\frac{1}{3}}$.∎